A Corporation uses a job-order costing system with a single plant-wide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on total fixed manufacturing overhead cost of $249,000, variable manufacturing overhead of $3.80 per machine-hour, and 30,000 machine-hours. The company has provided the following data concerning Job X784 which was recently completed:

Job E's manufacturing cost:

Total machine-hours

250

Direct materials

$

470

Direct labor cost

$

5,500

If the company marks up its unit product costs by 30% then the selling price for Job X784 is closest to:

Group of answer choices

A $2,698.50

B $12,693.50

C $7,761.00

D $11,693.50
Answer #1

Total variable overhead estimated=(3.8*30,000)=$114,000

Total overhead estimated=Total variable overhead estimated+Total fixed overhead estimated

=114,000+249,000

=$363000

Predetermined overhead rate=Total overhead estimated/Estimated machine hours

=$363000/30,000

=$12.1 per machine hour

Applied overhead=Predetermined overhead rate*Actual machine hours

=(12.1*250)=$3025

Add:Direct materials=$470

Add:Direct labor=$5500

Total cost=$8995

Hence total selling value=8995*1.3

=$11693.5